The guy wire on a TV antenna is 1 m longer than the height of the antenna. If the guy wire is anchored 3 m from the foot of the antenna, how tall is the antenna?

Answer by scott8148(6628)

by Pythagoras ___ x^2 + 3^2 = (x+1)^2

x^2 + 9 = x^2 + 2x + 1

8 = 2x

4 = x

a classic 3-4-5 triangle